A rigid vessel contains water at 400°F. Heat is to be added so that the water passes through the critical point. What should the quality be at the temperature of 400"F? Am. 0.01728
1
Expert's answer
2020-08-07T14:21:10-0400

Initial Temperature (T1)=400F=(5/9)(400-32)=204.44 C

If steam is heated at constant volume (in a rigid vessel) to teach critical point(Tc=374C),

T1=204.44C, T2=374C

From steam table :

At critical point specific volume(v2)=0.003105m3/kg

Therefore specific volume at initial state(v1)=(v2)=0.003105 m3/kg

At, T1=204.44C.

vf=0.001164 m3/kg, vg=0.11508 m3/kg

Quality of steam(X) at T1=204.44 C is given by

v1=vf+X(vg-vf)=0.001164+X(0.11508-0.001164 )=0.003105

X=0.01711

Ans: Quality of steam at T1=4000F is X=0.01711